Two straight paths are represented by equations x-3y=2 and -2x+6y=5. Check whether the path cross each other or not.

Grade:10

1 Answers

Rajendra Krishna
7 years ago
x -3y = 2
Multiply with -2
We get -2x + 6y = 6
Since the other line is -2x+6y =5 Both lines are parallel and never cross.
There will be no intersection